# Training-Free Label Space Alignment for Universal Domain Adaptation

### v1

@@ Abstract
Universal domain adaptation (UniDA) transfers knowledge from a labeled source domain to an unlabeled target domain, where label spaces may differ and the target domain may contain private classes. Previous UniDA methods primarily focused on visual space alignment but often struggled with visual ambiguities due to content differences, which limited their robustness and generalizability. To overcome this, we introduce a novel approach that leverages the strong zero-shot capabilities of recent vision-language foundation models (VLMs) like CLIP, concentrating solely on label space alignment to enhance adaptation stability. CLIP can generate task-specific classifiers based only on label names. However, adapting CLIP to UniDA is challenging because the label space is not fully known in advance. In this study, we first utilize generative vision-language models to identify unknown categories in the target domain. Noise and semantic ambiguities in the discovered labels—such as those similar to source labels (e.g., synonyms, hypernyms, hyponyms)—complicate label alignment. To address this, we propose a training-free label-space alignment method for UniDA (TLSA). Our method aligns label spaces instead of visual spaces by filtering and refining noisy labels between the domains. We then construct a universal classifier that integrates both shared knowledge and target-private class information, thereby improving generalizability under domain shifts. The results reveal that the proposed method considerably outperforms existing UniDA techniques across key DomainBed benchmarks, delivering an average improvement of +7.9%in H-score and +6.1% in H3-score. Furthermore, incorporating self-training further enhances performance and achieves an additional (+1.6%) increment in both H- and H3-scores.
@@ 1 Introduction
Domain shift remains a pivotal challenge, addressing the discrepancies between data distributions in real-world applications and training environments. Unsupervised domain adaptation (UDA) (Long et al., 2015; Ganin et al., 2016; Saito et al., 2018) aims to transfer knowledge from a labeled source domain to an unlabeled target domain, based on the stringent assumption that the label space remains unchanged across domains. However, in practical settings, the target label space may comprise only a subset of source classes that overlap with both domains or include target-exclusive classes unique to the target domain. Various studies have explored these complexities. Open-set domain adaptation (ODA) and partial-set domain adaptation (PDA) assume that the label space of each domain is a subset of another’s (Panareda Busto and Gall, 2017; Cao et al., 2018). Nonetheless, these approaches depend on prior knowledge of label-space shifts, limiting their real-world applicability. To address this constraint, Universal Domain Adaptation (UniDA) was introduced (You et al., 2019a; Saito et al., 2020) for broader applicability.
In UniDA, a primary challenge is distinguishing target-private samples from source-private ones. Numerous studies have proposed innovative detection criteria utilizing entropy (Saito et al., 2020; Qu et al., 2023a), binary classifiers (Saito and Saenko, 2021; Lu et al., 2024), optimal transport (Chang et al., 2022), energy-based methods (Chen et al., 2022a), class prototypes (Li et al., 2021a; Lai et al., 2023), and calibration techniques (Deng and Jia, 2023).Despite their capabilities, these methods struggle to differentiate between known and unknown classes that exhibit visual ambiguity due to content discrepancies across domains. This challenge arises because they rely on the geometric structure of visual features under domain shifts.
This study addresses these limitations by focusing exclusively on label-space alignment. To achieve this, we leverage the robust zero-shot capabilities of Vision-Language Models (VLMs) such as CLIP (Radford et al., 2021) and ALIGN (Li et al., 2021b). For instance, CLIP effectively recognizes diverse classes across domains by prompting label names through its fixed text encoder. To evaluate this ability, we conducted a pilot study assessing the performance of a zero-shot classifier across various label spaces within the UniDA framework. As shown in Fig. 2, zero-shot classifiers using customized label spaces exhibit substantial improvements in UniDA performance. Notably, simply incorporating oracle target private class names into the zero-shot classifier significantly outperformed state-of-the-art UniDA methods without requiring additional fine-tuning. These findings suggest that addressing the UniDA problem is more efficiently achieved by aligning label spaces rather than visual spaces across domains when utilizing VLMs. Moreover, the generated textual outputs are directly inspectable, enhancing the explainability and interpretability of the proposed pipeline.
Figure 1: Comparison between CLIP Prompt and state-of-the-art methods on H-score and H3-score. Our method achieves superior performance without additional training.
@@
Inspired by Fig. 2, we aim to identify target private classes and obtain an empirical version of these classes, termed predicted target private labels, by employing generative VLMs (Li et al., 2022, 2023). A similar approach is utilized in ODA (Zara et al., 2023), which discovers target private labels through text-to-text matching between the two domains at the cluster level. Nevertheless, this method is highly dependent on clustering quality. Moreover, texts identified from the source domain often lack informativeness due to domain shifts, particularly regarding abstract classes (Section 2.3). In contrast, we propose removing labels containing source-related information by conducting instance-level discovery exclusively on texts from the target domain.
@@
However, determining whether discovered labels include source-related information remains challenging due to noise (Fig. 2a) and semantic ambiguities, such as hyponyms (Fig. 2b) and synonyms (Fig. 2c)). For example, in Fig. 2b, the generative VLM outputs “Panasonic,” a hyponym of “Monitor” in the source labels. Without proper alignment, “Panasonic” is incorrectly predicted as a target private label, leading to degraded performance.
@@
To address this, we introduce Training-free Label Space Alignment for UniDA (TLSA). Specifically, we first identify the raw labels of the target samples using a generative VLM. Subsequently, we implement three steps to eliminate incorrect or semantically ambiguous discovered labels from the source labels and accurately predict the true target private labels: (i) Synonym label alignment (Section 3.3), which identifies synonyms in raw discovered labels using WordNet and aligns them with the source labels to remove synonyms; (ii) Semantic label alignment (Section 3.4). To resolve the ambiguous similarity between source and discovered labels, we identify a set of predictions that are considered correct for an image within the CLIP joint embedding space. When this set lacks source labels and contains only discovered labels, we select only the filtered discovered labels as candidate labels. (iii) Frequency-based noise-candidate filtering (Section 3.5). To eliminate residual noise, we monitor the frequency of target private candidate labels that pass previous filters across the entire target domain, removing noisy candidate labels based on their occurrence patterns. Ultimately, we obtain the predicted target private labels used to construct a universal classifier alongside the source labels. Our universal classifier more effectively differentiates between known and unknown labels than methods that depend on the geometric structure of visual representations (Chang et al., 2022; Lai et al., 2023; Wang et al., 2024a). Additionally, our model demonstrates further performance enhancements when integrated with self-training (Section  3.6).
@@
    We introduce a novel UniDA method by utilizing generative vision-language models to discover candidate target private classes through our filtering techniques.
@@
    By applying the filtered target private classes, we improve the CLIP zero-shot classifier for UniDA, termed the universal classifier.
@@
    Even without additional training, our approach achieves exceptionally strong performance compared to baselines.
@@
    We enhance the universal classifier through self-training to achieve full adaptation to the target domain.
  5. 5.
@@
Furthermore, as demonstrated in prior studies, fine-tuning the backbone CLIP on source data diminishes its ability to detect target private classes (Deng and Jia, 2023). This issue, known as feature distortion (Kumar et al., 2022; Shu et al., 2022; Trivedi et al., 2023; Saito et al., 2023), arises when the source data is used for tuning. Our method circumvents tuning on the source data by leveraging the source label space to filter and identify target private labels from the generative VLM. Consequently, our approach prevents feature distortion, enabling faster experimentation and improved performance.
@@ 2.1 Vision-Language Models
Figure 3: Comparison between existing UniDA methods and our proposed TLSA. Left: Prototype-based methods (Li et al., 2021a; Chang et al., 2022; Lai et al., 2023) construct target prototypes but struggle with visually. Right: Our TLSA identifies target private labels using only source labels and target samples.
@@ 2.2 Universal Domain Adaptation
Unlike UDA, UniDA assumes a distinct scenario where the label set of the target domain may partially intersect with that of the source domain. Determining whether a given data belongs to a known or unknown class is a highly challenging task because the model can only predict the classes it is trained on. Therefore, baseline methods for UDA, such as adversarial learning (Ganin et al., 2016; Tzeng et al., 2017; Long et al., 2018) and metric learning (Long et al., 2015, 2017), lead to degradation in performance. The problem occurs because domain alignment is attempted without considering the possibility that the label space might differ, leading to negative transfer.
With the recent progress of VLMs, there has been growing recognition of the relatively limited exploration compared to pretrained backbones on ImageNet (Russakovsky et al., 2015) such as ResNet (He et al., 2016) and ViT (Dosovitskiy et al., 2020). In response, CLIP-Distillation (Deng and Jia, 2023) was proposed, utilizing foundation models (Radford et al., 2021; Oquab et al., 2023) trained on large datasets through self-supervised learning. This approach achieved state-of-the-art performance among CLIP-pretrained backbones by performing auto-calibration via learning the classifier temperature. Similarly, MemSPM (Lai et al., 2023) adopts a memory-based approach that models sub-classes by storing and updating sub-prototypes, while keeping the CLIP backbone frozen. Here, CLIP is employed not for its zero-shot capability, but merely as a generic feature extractor. Fig. 3 illustrates the key differences between prior works and our approach. Unlike previous methods, and as motivated by Fig. 2, we reformulate UniDA as the simpler task of discovering target-private class names with generative VLMs. Specifically, we generate candidate labels directly from target-domain samples using a generative VLM, and then apply a filtering strategy that exploits image–text similarity to refine them.
@@ 2.3 Out-Of-Distribution Detection
Among these, the recent study aims to leverage CLIP’s zero-shot capability for open-set domain adaptation (Zara et al., 2023). However, there are significant limitations when applying this method (Zara et al., 2023) directly to UniDA. The number of clusters, K, is determined based on prior knowledge from the open-set, where the number of clusters exceeds the number of source labels. In a universal set, the decision of K is more challenging as its extent is not pre-defined. Specifically, when using K-means to predict the number of classes, the elbow method is typically applied (Hartigan and Wong, 1979). However, the elbow point is often unreliable due to a smooth score plot, leading to inaccuracies. Furthermore, we argue that texts discovered from the source domain are ineffective for detecting target private classes due to domain shifts, as the shapes of objects can differ entirely between the source and target domains, especially for abstract labels.
@@ 2.4 Self-Training
Self-training is a widely studied approach in semi-supervised learning that has recently been explored for enhancing the performance of minority classes (Amini et al., 2024). In situations where domain shift exists, research has primarily focused on Source-free domain adaptation (SFDA) and Test time adaptation (TTA). One of the most well-known approaches is a pseudo labeling technique, which is both simple to implement and intuitive. Pseudo labeling assumes that predictions for confident samples are trustworthy. Initially, training leverages these trustworthy predictions, and as the model advances, the scope of pseudo labeling is progressively broadened. The most important aspect is to maintain class balance. Class imbalance can lead to overconfidence in certain classes, which ultimately degrades the model’s performance. To mitigate this issue, various source-free domain adaptation methods, such as SHOT (Liang et al., 2020) and DIFO (Tang et al., 2024) incorporate a diversity term to perform entropy maximization. ST-SGG (Kim et al., 2024) devises a method to adapt class-adaptive thresholds based on frequency, adjusting them according to the situation in previous iterations.
Figure 4: Overview of TLSA. We first gather candidate target-private labels from a generative VLM, then refine them to train a universal classifier. (a) Synonym alignment (Section 3.3) removes source label synonyms via WordNet: (b) Semantic alignment (Section 3.4) re-scores image-label pairs in the embedding space and decides shared vs. target-private; (c) Frequency filtering (Section 3.5) prunes noisy low support candidates using frequency banks; (d) Self-training (Section 3.6) applies a teacher-student scheme on class-balanced top-k confident samples.
Inspired by this research direction, we choose to select the top-k most confident samples for each predicted class as pseudo labels to ensure class balance. Furthermore, to maintain this balance among the pseudo labels, we define k as the batch size divided by the number of predicted classes. This approach limits the selection to samples that can be easily classified. Additionally, we empirically confirm that our universal classifier, created by predicted target private labels, discriminates better than the $|C_{s}|$+1 way classifier, where $|C_{s}|$ is the number of source classes (Fig. 7).
@@ 3.1 Problem Definition
In UniDA, there is a labeled source domain $\mathcal{D}^{s}=\{(x^{s}_{i},y^{s}_{i})\}^{N_{s}}_{i=1}$ where $x^{s}_{i}\in\mathcal{X}^{s},y^{s}_{i}\in\mathcal{Y}^{s}\subset\mathbb{R}^{|C_{s}|}$, and an unlabeled target domain $\mathcal{D}^{t}=\{x^{t}_{j}\}^{N_{t}}_{j=1}$ where $x^{t}_{j}\in\mathcal{X}^{t}$. We denote the ground-truth labels of the target domain as $\mathcal{Y}^{t}\subset\mathbb{R}^{|C_{t}|}$. $|C_{s}|$ represents the number of source labels, and $|C_{t}|$ represents the number of target labels. $C=C_{s}\cap C_{t}$ denotes the common label sets shared by source and target domains. $\bar{C_{s}}=C_{s}-C$ and $\bar{C_{t}}=C_{t}-C$ stand for source-private label and target private label, respectively. Our goal is to filter the raw discovered label $R$ from the generative VLM to align the ground-truth target private classes $\bar{C_{t}}$.

### v2

@@ Abstract
Universal domain adaptation (UniDA) transfers knowledge from a labeled source domain to an unlabeled target domain, where label spaces often differ and the target domain often contains private classes. Previous UniDA methods primarily focused on visual space alignment, however, severe domain shifts often lead to visual ambiguities, undermining their ability to predict the label space shift. Our preliminary experiment shows that focusing exclusively on label space alignment can be a more effective solution to these limitations. While generative vision-language models (VLMs) offer strong zero-shot capabilities that could help uncover target-private classes, naively employing them introduces challenges such as noisy and semantically ambiguous labels. To address these challenges, we propose a novel approach that carefully utilizes generative VLMs to improve label space alignment. We introduce an adaptive thresholding strategy that leverages meaningful relationships between source and discovered target labels. This allows a training-free mechanism to filter out noisy and ambiguous labels, effectively refining the label space by identifying shared and discovering target-private label subsets. We then build a universal classifier for robust self-training, which effectively utilizes reliable pseudo-labels derived from the identified label spaces. The results reveal that the proposed method considerably outperforms existing UniDA techniques across key DomainBed benchmarks, delivering an average improvement of +7.9% in H-score and +6.1% in H3-score. Furthermore, incorporating self-training further enhances performance and achieves an additional (+1.6%) increment in both H- and H3-scores.
@@
[…]
[inst1]organization=Department of Artificial Intelligence, Korea University,addressline=145 Anam-ro, Seongbuk-gu, city=Seoul, postcode=02841, country=South Korea
[…]
[inst2]organization=OMRON SINIC X,addressline=5-24-5 Hongo, Bunkyo-ku, city=Tokyo, postcode=113-0033, country=Japan
@@ 1 Introduction
Domain shift, the discrepancy between training and real-world data, remains a pivotal challenge for learning generalizable vision-language representations. Unsupervised domain adaptation (UDA) (Long et al., 2015; Ganin et al., 2016; Saito et al., 2018) aims to transfer knowledge from a labeled source domain to an unlabeled target domain, based on the stringent assumption that the label space remains unchanged across domains. However, in practical settings, the target label space often comprises only a subset of source classes that overlap with both domains or include target-exclusive classes unique to the target domain. Various studies have explored these complexities. Open-set domain adaptation (ODA) and partial-set domain adaptation (PDA) assume that the label space of each domain is a subset of another’s (Panareda Busto and Gall, 2017; Cao et al., 2018). Nonetheless, these approaches rely on prior knowledge of label space shifts, which limits their real-world applicability. To address this constraint, Universal Domain Adaptation (UniDA) was introduced (You et al., 2019a; Saito et al., 2020) for broader applicability.
Figure 1: Comparison between CLIP Prompt, a zero-shot classifier enhanced with corresponding label text prompts, and previous state-of-the-art (SOTA) methods. Simply adding target-private classes or removing source-private classes leads to substantial gains in H-score, achieved entirely without additional training.
@@
In UniDA, a primary challenge is distinguishing target-private samples from source-private ones. Many studies have proposed innovative detection criteria utilizing entropy (Saito et al., 2020; Qu et al., 2023a), binary classifiers (Saito and Saenko, 2021; Lu et al., 2024), optimal transport (Chang et al., 2022), energy-based methods (Chen et al., 2022a), class prototypes (Li et al., 2021a; Lai et al., 2023), and calibration techniques (Deng and Jia, 2023). These methods still suffer from severe domain shifts, which involve not only style variations but also content changes among shared classes. The reason is that the methods rely on the geometric structure of visual embeddings; however, such features are often drastically shifted across domains, leading to inconsistent decisions.
This study addresses these limitations by focusing exclusively on label space alignment. To achieve this, we leverage the robust zero-shot capabilities of Vision-Language Models (VLMs) such as CLIP (Radford et al., 2021) and ALIGN (Li et al., 2021b). For instance, CLIP effectively recognizes diverse classes across domains by prompting label names through its fixed text encoder. To evaluate this ability, we conducted a pilot study assessing the performance of a zero-shot classifier across various category spaces within the UniDA framework. As shown in Fig. 2, zero-shot classifiers using customized label spaces exhibit substantial improvements in UniDA performance. Interestingly, merely augmenting the zero-shot classifier with oracle category spaces already surpasses existing UniDA approaches, most of which attempt to achieve alignment solely in the visual space. These findings suggest that addressing the UniDA problem is more efficiently achieved by aligning label spaces rather than visual spaces across domains.
@@
Inspired by Fig. 2, we aim to identify target-private classes and obtain an empirical version of these classes, termed predicted target-private labels, by employing generative VLMs (Li et al., 2022, 2023). In particular, predicted target-private labels are derived from labels discovered in target images using generative VLMs. Nevertheless, distinguishing whether discovered labels belong to part of the source labels or not becomes challenging due to noise (Fig. 2a) and semantic ambiguities, such as hyponyms (Fig. 2b) and synonyms (Fig. 2c). For example, in Fig. 2b, the generative VLM outputs “Panasonic,” a semantic hyponym of “Monitor” in the source labels. Without proper alignment, “Panasonic” is incorrectly predicted as a target-private label, leading to degraded performance.
@@
To address this, we introduce Training-free Label Space Alignment for UniDA (TLSA). Specifically, we first identify the raw labels of the target samples using a generative VLM. Subsequently, we implement three steps to eliminate incorrect or semantically ambiguous discovered labels from the source labels and accurately predict the true target private labels: (i) Synonym label alignment (Sec. 3.3): We identify synonyms in raw discovered labels using WordNet and align them with the source labels to remove synonyms; (ii) Semantic label alignment (Sec. 3.4): To resolve the ambiguous similarity between source and discovered labels, we identify a set of predictions that are considered correct for an image within the CLIP joint embedding space through an adaptive thresholding process. If the resulting set excludes source labels and includes only discovered labels, the filtered discovered labels are retained as candidate labels; (iii) Frequency-based noise-candidate filtering (Sec. 3.5): To eliminate residual noise, we monitor the frequency of target private candidate labels that pass previous filters across the entire target domain, removing noisy candidate labels based on their occurrence patterns.
Lastly, leveraging TSLA, we derive the predicted target-private labels, which are used together with source labels to construct a universal classifier for robust self-training. Compared to methods that rely on the geometric structure of visual representations (Chang et al., 2022; Lai et al., 2023; Wang et al., 2024a), our universal classifier more effectively distinguishes between known and unknown labels and produces more reliable pseudo-labels, leading to further performance enhancements (Sec. 3.6).
@@
    We propose a novel training-free label space alignment method for Universal Domain Adaptation (UniDA), shifting the focus from visual feature alignment to label space alignment using generative vision-language models (VLMs).
@@
    To address semantic challenges in naïve label space alignment with generative VLMs, we propose new filtering techniques to resolve semantic ambiguities and noise, yielding a refined shared label space between source and target domains, as well as a well-defined target-private label space.
@@
    By applying the filtered target private classes, we enhance the CLIP zero-shot classifier for UniDA, referred to as the universal classifier, which can be further enhanced through self-training to achieve full adaptation to the target domain.
@@
Furthermore, as demonstrated in prior studies, fine-tuning the backbone CLIP on source data diminishes its ability to detect target private classes (Deng and Jia, 2023). This issue, known as feature distortion (Kumar et al., 2022; Shu et al., 2022; Trivedi et al., 2023; Saito et al., 2023), arises when the source data is used for tuning. Our method circumvents tuning on the source data by leveraging the source label space to filter and identify target private labels from the generative VLM. Furthermore, our approach prevents feature distortion, enabling faster experimentation and improved performance.
Figure 3: Comparison between prior UniDA methods and our TLSA+ST. Left: Visual space alignment approaches (Li et al., 2021a; Chang et al., 2022; Lai et al., 2023) rely on cluster matching, which becomes ambiguous under domain shift. Right: Our TLSA+ST discovers labels from target images and aligns them with source labels with adaptive thresholding using the distance in the embedding space between source labels. Since TLSA+ST uses pretrained CLIP’s joint embedding space, it is robust against domain shift.
@@ 2.2 Universal Domain Adaptation
Unlike UDA, UniDA assumes a distinct scenario where the label set of the target domain often partially intersects with that of the source domain. Determining whether a given data belongs to a known or unknown class is a highly challenging task because the model can only predict the classes it is trained on. Therefore, baseline methods for UDA, such as adversarial learning (Ganin et al., 2016; Tzeng et al., 2017; Long et al., 2018) and metric learning (Long et al., 2015, 2017), lead to degradation in performance. The problem occurs because domain alignment is attempted without considering the possibility that the label space might differ, leading to negative transfer.
With the recent progress of VLMs, there has been growing recognition of the relatively limited exploration compared to pretrained backbones on ImageNet (Russakovsky et al., 2015), such as ResNet (He et al., 2016) and ViT (Dosovitskiy et al., 2020). In response, CLIP-Distillation (Deng and Jia, 2023) was proposed, utilizing foundation models (Radford et al., 2021; Oquab et al., 2023) trained on large datasets through self-supervised learning. This approach achieved state-of-the-art performance among CLIP-pretrained backbones by performing auto-calibration via learning the classifier temperature. Similarly, MemSPM (Lai et al., 2023) adopts a memory-based approach that models sub-classes by storing and updating sub-prototypes, while keeping the CLIP backbone frozen. Here, CLIP is employed not for its zero-shot capability, but merely as a generic feature extractor. Fig. 3 illustrates the key differences between prior works and our approach. Unlike previous methods, and as motivated by Fig. 2, we reformulate UniDA as the simpler task of discovering target-private class names. Specifically, we generate candidate labels directly from target-domain samples using a generative VLM, and then apply a filtering strategy that exploits image–text similarity to refine them.
@@ 2.3 Out-Of-Distribution Detection
Among these, the recent study aims to leverage CLIP’s zero-shot capability for open-set domain adaptation (Zara et al., 2023). However, there are significant limitations when applying this method directly to UniDA. The number of clusters, $k$, is determined based on prior knowledge from the open-set, where the number of clusters exceeds the number of source labels. In a universal set, the decision of $k$ is more challenging as its extent is not pre-defined. Specifically, when using K-means to predict the number of classes, the elbow method is typically applied (Hartigan and Wong, 1979). However, the elbow point is often unreliable due to a smooth score plot, leading to inaccuracies. Furthermore, we argue that texts discovered from the source domain are ineffective for detecting target private classes due to domain shifts, as the shapes of objects can differ entirely between the source and target domains, especially for abstract labels.
@@ 2.4 Self-Training
Self-training is a widely studied approach in semi-supervised learning that has recently been explored for enhancing the performance of minority classes (Amini et al., 2024). In situations where domain shift exists, research has primarily focused on Source-free domain adaptation (SFDA) and test-time adaptation (TTA). One of the most well-known approaches is a pseudo labeling technique, which is both simple to implement and intuitive. Pseudo labeling assumes that predictions for confident samples are trustworthy. Initially, training leverages these trustworthy predictions, and as the model advances, the scope of pseudo labeling is progressively broadened. The most important aspect is to maintain class balance. Class imbalance can lead to overconfidence in certain classes, which ultimately degrades the model’s performance. To mitigate this issue, various source-free domain adaptation methods, such as SHOT (Liang et al., 2020) and DIFO (Tang et al., 2024), incorporate a diversity term to perform entropy maximization. ST-SGG (Kim et al., 2024) devises a method to adapt class-adaptive thresholds based on frequency, adjusting them according to the situation in previous iterations.
Inspired by this research direction, we choose to select the top-k most confident samples for each predicted class as pseudo labels to ensure class balance. Furthermore, to maintain this balance among the pseudo labels, we define $k$ as the batch size divided by the number of predicted classes. This approach limits the selection to samples that can be easily classified. Additionally, we empirically confirm that our universal classifier, created by predicting target private labels, discriminates better than the $|C_{s}|+1$ way classifier, where $|C_{s}|$ is the number of source classes (see Fig. 8).
@@ 3.1 Problem Definition
In UniDA, we are given a labeled source domain $\mathcal{D}^{s}=\{(x^{s}_{i},y^{s}_{i})\}^{N_{s}}_{i=1}$ where $x^{s}_{i}\in\mathcal{X}^{s},y^{s}_{i}\in\mathcal{Y}^{s}\subset\mathbb{R}^{|C_{s}|}$, and an unlabeled target domain $\mathcal{D}^{t}=\{x^{t}_{j}\}^{N_{t}}_{j=1}$ where $x^{t}_{j}\in\mathcal{X}^{t}$. We denote the ground-truth labels of the target domain as $\mathcal{Y}^{t}\subset\mathbb{R}^{|C_{t}|}$. $|C_{s}|$ represents the number of source labels, and $|C_{t}|$ represents the number of target labels. $C=C_{s}\cap C_{t}$ denotes the common label sets shared by source and target domains. $\bar{C_{s}}=C_{s}-C$ and $\bar{C_{t}}=C_{t}-C$ stand for source-private label and target private label, respectively. Our goal is to filter the raw discovered label $R$ from the generative VLM to align the ground-truth target private classes $\bar{C_{t}}$.
Figure 4: Overview of TLSA. We first gather candidate target-private labels from a generative VLM, then refine them to train a universal classifier. (a) Synonym label alignment (Sec. 3.3) removes source label synonyms via WordNet: (b) Semantic label alignment (Sec. 3.4) re-scores image-label pairs in the embedding space and decides shared vs. target-private; (c) Frequency-based noisy candidate filtering (Sec. 3.5) prunes noisy low support candidates using frequency banks; and (d) Self-training (Sec. 3.6) applies a teacher-student scheme on class-balanced top-k confident samples.